OpenStudy (lena772):

A pair of linear equations is shown below: y = –x + 1 y = 2x + 4 Which of the following statements best explains the steps to solve the pair of equations graphically? On a graph, plot the line y = –x + 1 which has y intercept = –1 and slope = 1; and y = 2x + 4 which has y intercept = 2 and slope = 4; and write the coordinates of the point of intersection of the two lines as the solution.

OpenStudy (lena772):

On a graph, plot the line y = –x + 1 which has y intercept = 1 and slope = 1; and y = 2x + 4 which has y intercept = 1 and slope = 4; and write the coordinates of the point of intersection of the two lines as the solution. On a graph, plot the line y = –x + 1 which has y intercept = 1 and slope = –1; and y = 2x + 4 which has y intercept = –2 and slope = 2; and write the coordinates of the point of intersection of the two lines as the solution. On a graph, plot the line y = –x + 1 which has y intercept = 1 and slope = –1; and y = 2x + 4 which has y intercept = 4 and slope = 2; and write the coordinates of the point of intersection of the two lines as the solution.

OpenStudy (lena772):

Yw